Stock X has a 9.5% expected return, a beta coefficient of 0.8, and a 30% standard deviation of expected returns. Stock Y hasHi there! I need help with A, C, and E, please. Thanks :)
Answer #1

A. Coefficient of variation of X=Standard Deviation/Expected Return =30%/9.5%=3.16
Coefficient of variation of Y =Standard Deviation/Expected Return =30%/12.5%=2.4

C Required Rate of X =Risk free rate+Beta*(Market Return -Risk Free rate) =6%+0.8*5% =10%
Required Rate of Y =Risk free rate+Beta*(Market Return -Risk Free rate) =6%+1.2*5% =12%

E. Required rate of return of portfolio =Weight of X*Return of X+weight of Y*Return of Y
=5500/7500*10%+2000/7500*12% =10.53%